# Automatic infant 2D pose estimation from videos: comparing seven deep neural network methods

### v2

@@ Abstract
Automatic markerless estimation of infant posture and motion from ordinary videos carries great potential for movement studies “in the wild”, facilitating understanding of motor development and massively increasing the chances of early diagnosis of disorders. There is rapid development of human pose estimation methods in computer vision thanks to advances in deep learning and machine learning. However, these methods are trained on datasets featuring adults in different contexts. This work tests and compares seven popular methods (AlphaPose, DeepLabCut/DeeperCut, Detectron2, HRNet, MediaPipe/BlazePose, OpenPose, and ViTPose) on videos of infants in supine position. Surprisingly, all methods except DeepLabCut and MediaPipe have competitive performance without additional finetuning, with ViTPose performing best. Next to standard performance metrics (object keypoint similarity, average precision and recall), we introduce errors expressed in the neck-mid-hip ratio and additionally study missed and redundant detections and the reliability of the internal confidence ratings of the different methods, which are relevant for downstream tasks. Among the networks with competitive performance, only AlphaPose could run close to real time (27 fps) on our machine. We provide documented Docker containers or instructions for all the methods we used, our analysis scripts, and processed data at https://hub.docker.com/u/humanoidsctu and https://osf.io/x465b/.
Introduction
@@
Accurate motion data are key to understand how infants develop and to identify deviations from normal development. Trained experts can identify the risks of developmental disorders, such as cerebral palsy, from spontaneous movements (using the General Movement Assessment (GMA) [1]) or through a neurological examination based on movement, posture, and reflexes (Hammersmith Infant Neurological Examination (HINE) [2]). However, expert evaluation constitutes a critical bottleneck in the early detection of developmental disorders, especially in less developed countries.
@@
Automated extraction and evaluation of movement patterns constitutes a key enabling technology to make screening available to a much larger population. In general, there are two classes of methods for collecting motion data [3]: (i) direct sensing where movements are captured using hardware attached to the bodies (e.g., inertial sensors and magnetic tracking systems) and (ii) indirect sensing (e.g., 3D motion capture, RGB cameras, RGB-D cameras). From the first group (see [4] for a review of wearable sensor systems to monitor body movements of neonates), the most popular are inertial sensors. These can be accelerometers [5], or, more frequently, inertial measurement units (IMUs – 3-axis accelerometer, 3-axis gyroscope, 3 magnetometers) integrated into a wearable suit (e.g., [6]). Wearable inertial sensors may be relatively inexpensive (compared to 3D motion capture), but the fact that they need to be physically attached to infants prevents their widespread use and may also affect spontaneous movement production. Moreover, inertial sensors do not provide absolute position information but only angular velocities and linear accelerations, which limits the analyses that can be run on the data (position information can be obtained from numerical integration but is prone to noise). Within indirect sensing, we want to draw a line between marker-based 3D motion capture systems and standard video cameras (RGB). On one end of the spectrum, 3D motion capture systems typically use infrared retroreflective markers placed on infant bodies and record their positions with multiple cameras. This can yield submillimeter accuracy but requires very expensive equipment and the markers may affect spontaneous movement production. Furthermore, markers cannot be placed on the back of infants in supine position which prevents the use of standard skeleton solvers. Markers may also be used with standard video cameras [7]. Multi-camera setups (e.g., 9 high-frequency cameras with studio lighting in [8]) provide an alternative to marker-based motion capture systems. On the other end of the spectrum are standard video cameras (e.g., [9]), like those in consumer cell phones, which constitute the only device truly accessible to almost everyone, without any barriers. Recordings from two cameras or RGB-D cameras containing depth information (e.g., Microsoft Kinect, Intel Realsense) could still be relatively easily applied “in the wild”.
@@
The next step after motion data acquisition is their analysis. The key application is automated clinical movement assessment. The methods can be classified based on the type of input data (e.g., accurate absolute 3D positions of body parts from a motion capture on one end of the spectrum or a single video stream from an ordinary camera on the other end) and on the assessment being performed—see [10] for a review. Most often, the desired output is automated GMA classification (see [11, 12] for reviews), with risks of cerebral palsy prediction as the typical clinical outcome. Direct sensing methods (e.g., accelerometers in [5]) or 3D data from marker-based motion capture have been used (e.g., [13]) but our focus is on indirect sensing and video-based approaches in particular (see Silva et al.[12] for a review). The images of infants can be used to directly extract features that are used for subsequent analyzes and assessment (e.g., [14] for GMA or [15] analyzing U-shaped developmental changes). With the rapid advent of pose estimation methods in computer vision and machine learning, methods that use 2D keypoint extraction in infant images are becoming increasingly popular. McCay et al. [16] extracted the positions of infant body keypoints using OpenPose [17], computed histograms of joint orientation and displacement in 2D and then used deep learning to classify abnormal movements (GMA labels). Chambers et al. [18] used the keypoints obtained from OpenPose as input to extract movement features and then applied a Naïve Bayesian Suprise Metric to predict neuromotor development risk. Reich et al. [19] used the positions of the 25 keypoints obtained from OpenPose directly for classification and evaluated the agreement with GMA. Shin et al. [9] obtained the positions of 2D keypoints using AlphaPose [20], estimated joint angle values from the 2D skeletons, analyzed complexity and then applied deep learning for classification (reporting correlations with HINE). Similarly, the risk of autism spectrum disorder (ASD) can be predicted from infant video recordings—using input images directly [21] or using 2D pose estimation (OpenPose) first [22].
Next to automated clinical assessment, movement analysis of infants is key to understanding normal development. Insights from spontaneous infant behavior often draw on small datasets and simplified but laborious manual scoring of video recordings (e.g., [23]). Quantitative data such as kinematics from motion capture are an exception (e.g., [24, 25]) and the length of the recordings is limited. Daily spontaneous recordings of infants will be needed [26] to uncover developmental trajectories on multiple and nested time scales and to employ nonlinear or dynamical systems analysis tools (e.g., [24]).
Sharing datasets including raw video footage is critical to make progress in psychological science [27, 28]. However, automatic and accurate extraction of motion data from this material is a critical prerequisite for progress. This is where this article ties in.
@@
Methods for automatic human keypoint extraction (for example, eyes, wrists, hips, feet, etc.), also called human pose estimation, from images and videos are rapidly evolving, with performance increasing every year (see [29, 30, 31] for recent surveys). These provide a key-enabling technology to make automatic motion extraction and analysis from “in the wild” recordings possible. These methods were developed primarily to detect and extract postures from adult bodies. So far, these methods have been applied to infant videos “as-they-are”. However, the morphology of the infant body is different from the body proportions of the adult, especially in early infancy [32, 33]. Additionally, small infants are typically in supine position (on their back) and move differently than standing adults, which constitutes a dataset which is out of the distribution the adult pose estimation models were trained on. Although automated pose estimation seems to be generally good (see Fig. 1 a-c), in a previous study [34], we encountered some limitations of current pose estimation methods in infants. For example, OpenPose [17] was observed to struggle with some camera angles, body postures, and keypoints. In particular, complex leg positions, such as when legs are crossed in supine position, are not often estimated with accuracy. This can be seen in Fig. 1 (d): OpenPose misses the left ankle keypoint and places the left knee keypoint on the thigh, very close to the left hip. In the current study, comparing different pose estimation methods, we found similar examples. Fig. 1 (e) shows HRNet Bottom-Up failing on a complex leg position and, maybe as a consequence of the messed leg estimation, failing to place the left shoulder and elbow keypoints, placing them on the right; while (c) and (f) show that HRNet Top-Down and Detectron2 detect a second person in the image where there is none or with highly misplaced keypoints.
@@
Figure 1: Examples of successful and erroneous keypoint estimations obtained from the different human pose estimation methods: a) ViTPose (infant TH, 17 weeks), b) HRnet Bottom-Up (infant AA, 17 weeks), c) HRnet Top-Down (TH, 10weeks, st2) with a successful estimation but a second hallucinated person, d) OpenPose missing the left ankle keypoint (synthetic infant 1); e) HRNet Bottom-Up failing to estimate a complex leg position (synthetic infant 10); f) Detectron2 detecting an additional infant with misplaced keypoints (synthetic infant 12).
@@
Thus, our first goal and contribution is to assess the performance of 2D human pose estimation methods when applied specifically to infants in supine position. With the exception of Needham et al. [8] who have compared OpenPose, AlphaPose, and DeepLabCut, this has not been systematically studied. In this article, we compare the performance of seven methods: Alphapose [20], DeepLabCut / DeeperCut (v2.2.1.1) [35], Detectron2 (v0.6) [36], MediaPipe / BlazePose (v0.10.14) [37, 38], HRNet (mmpose v0.28.0) [39, 40] (Bottom-Up (HRnet BU) and Top-Down (HRnet TD)), OpenPose [17] and ViTPose (mmpose v1.1.0) [41].
@@
The third contribution of this work constitutes in making all the pose estimation methods and tools for the results analysis available to the community. Although the pose estimation algorithms themselves are publicly accessible in their corresponding repositories, we make available the versions we used including the complete environment in public Docker containers (at https://hub.docker.com/u/humanoidsctu). In addition, we also share the evaluation scripts and the detailed results (at https://osf.io/x465b/).
@@ Methods
To ensure that comparisons are as fair as possible, we used versions of the pose estimation methods that were trained with the same dataset: COCO [42], although some methods only have available weights that used other datasets in their training, instead of or in addition to COCO (see each method’s respective subsection below). The methods provide 17 estimated keypoints, with the exception of DeepLabCut, MediaPipe and OpenPose that provide 14, 33 and 18 keypoints respectively. We will call a single set of estimated keypoints a detection. The details of the parameters are provided in the Supplementary Materials. All methods except MediaPipe provide a form of confidence values as their own internal evaluation of the estimated result, with higher values indicating higher confidence in the quality of the estimation. This value can be provided for the whole set of keypoints or for individual keypoints (or both). Following the trend in the documentation and outputs of the methods and how they use these words, we will use score when referring to the whole detection value and confidence when referring to the individual keypoints values. The methods follow either of the two approaches: Top-Down, which first uses a detector to delimit areas of interest in the image that encompass a person to detect, and then estimates a set of keypoints within this space; and Bottom-Up, which finds keypoints or sets of keypoints separately on the whole image before joining them into full keypoints sets and separating them into several individuals if necessary. A summary of the methods’ properties and description is available in the Supplementary Materials Table ST.1.
AlphaPose
@@
DeepLabCut/DeeperCut
We use the DeepLabCut environment [43], with its human pose estimation DeeperCut [35] based on an R-CNN architecture. It follows a Bottom-Up approach. It has been trained on the MPII dataset [44] and provides 14 keypoints. It provides confidences but no scores.
Detectron2
@@
MediaPipe/BlazePose
While commonly known as MediaPipe [37], this environment’s human pose estimation uses BlazePose [38] and is based on a CNN architecture, following a Top-Down approach. It has been trained on a custom dataset with additional face and hand annotations on top of COCO’s regular 17 keypoints. It provides neither scores nor confidences.
HRnet Bottom-Up and Top-Down
HRNet [39, 40] is named after the name of the neural architecture it uses, a High Resolution Network, based on CNNs. It has been trained using both COCO and MPII datasets. We use its implementation through the MMPose environment [45], which proposes bottom-up (BU) and top-down (TD) versions of HRNet. HRNet BU provides scores, while HRNet TD provides confidences.
OpenPose

### v3

@@ Abstract
Automatic markerless estimation of infant posture and motion from ordinary videos carries great potential for movement studies ”in the wild”, facilitating understanding of motor development and massively increasing the chances of early diagnosis of disorders. There is rapid development of human pose estimation methods in computer vision thanks to advances in deep learning and machine learning. However, these methods are trained on datasets that feature adults in different contexts. This work tests and compares seven popular methods (AlphaPose, DeepLabCut/DeeperCut, Detectron2, HRNet, MediaPipe/BlazePose, OpenPose, and ViTPose) on videos of infants in supine position and in more complex settings. Surprisingly, all methods except DeepLabCut and MediaPipe have competitive performance without additional finetuning, with ViTPose performing best. Next to standard performance metrics (average precision and recall), we introduce errors expressed in the neck-mid-hip (torso length) ratio and additionally study missed and redundant detections, and the reliability of the internal confidence ratings of the different methods, which are relevant for downstream tasks. Among the networks with competitive performance, only AlphaPose could run close to real time (27 fps) on our machine. We provide documented Docker containers or instructions for all the methods we used, our analysis scripts, and the processed data at https://hub.docker.com/u/humanoidsctu and https://osf.io/x465b/.
1 Introduction
Movement analysis of infants is key to understanding their development. Insights from spontaneous infant behavior often draw on small datasets and simplified but laborious manual scoring of video recordings (e.g., [1]). Quantitative data such as kinematics from motion capture are an exception [2, 3], and the length of the recordings is limited. Daily and longer spontaneous recordings of infants will be needed [4] to uncover developmental trajectories on multiple and nested time scales and to employ nonlinear or dynamical systems analysis tools [2]. The sharing of data sets, including raw video footage, is critical to making progress in psychological science [5, 6]. Some efforts in this direction are under way, but manually scoring hours, days, and weeks of recordings becomes a bottleneck. Automatic and accurate extraction of motion data from this material is a critical prerequisite for progress. One example is the analysis of kinematics during reaching development. In order to obtain the relevant features (e.g., average speed, movement duration, maximum speed, jerk, path length, distance, straightness ratio) [7, 8] infants need to be tested in the laboratory, and motion capture markers or similar equipment is typically used, which limits the amount of data that can be collected and may even affect the infants’ performance on the task. Another example are infants’ self-directed spontaneous movements like self-touch. In [9, 1, 10], they were scored manually, which is extremely laborious. Both manual and automatic scoring has been used in [11]. The key benefits of automatic pose estimation are thus the following: (i) recordings can be performed in the wild—at homes and without additional instrumentation; (ii) very long recordings can be automatically analyzed and reliability can be established by manually scoring only a subset of the data; (iii) complete 2D postures and movement trajectories are available.
@@ 1 Introduction
Another key area is clinical practice—identifying deviations from normal development. Trained experts can identify the risks of developmental disorders, such as cerebral palsy, from spontaneous movements (using the General Movement Assessment (GMA) [12]) or through a neurological examination based on movement, posture, and reflexes (Hammersmith Infant Neurological Examination (HINE) [13]). However, expert evaluation constitutes a critical bottleneck in the early detection of developmental disorders, especially in less developed countries. Automated extraction and evaluation of movement patterns constitutes a key enabling technology to make screening available to a much larger population.
@@
In general, there are two classes of methods for collecting motion data [14]: (i) direct sensing where movements are captured using hardware attached to the bodies (e.g., inertial sensors and magnetic tracking systems) and (ii) indirect sensing (e.g., 3D motion capture, RGB cameras, RGB-D cameras). From the first group (see [15] for a review of wearable sensor systems to monitor body movements of neonates), the most popular are inertial sensors. These can be accelerometers [16], or, more frequently, inertial measurement units (IMUs – 3-axis accelerometer, 3-axis gyroscope, 3 magnetometers) integrated into a wearable suit [17]. Wearable inertial sensors may be relatively inexpensive (compared to 3D motion capture), but the fact that they need to be physically attached to infants prevents their widespread use and may also affect spontaneous movement production. Moreover, inertial sensors do not provide absolute position information but only angular velocities and linear accelerations, which limits the analyses that can be run on the data (position information can be obtained from numerical integration but is prone to noise). Within indirect sensing, we want to draw a line between marker-based 3D motion capture systems and standard video cameras (RGB). On one end of the spectrum, 3D motion capture systems typically use infrared retroreflective markers placed on infant bodies and record their positions with multiple cameras. This can yield sub-millimeter accuracy but requires very expensive equipment, and the markers may affect spontaneous movement production. Furthermore, markers cannot be placed on the back of infants in supine position which prevents the use of standard skeleton solvers. Markers may also be used with standard video cameras [18]. Multi-camera setups (e.g., 9 high-frequency cameras with studio lighting in [19]) provide an alternative to marker-based motion capture systems. On the other end of the spectrum are standard video cameras [20], such as those in consumer cell phones, which constitute the only device that is truly accessible to almost everyone, without any barriers. Recordings from two cameras or RGB-D cameras containing depth information (e.g., Microsoft Kinect, Intel Realsense) could still be relatively easily applied “in the wild”.
@@
The next step after motion data acquisition is their analysis. One of the key applications is automated clinical movement assessment. The methods can be classified based on the type of input data (e.g., accurate absolute 3D positions of body parts from a motion capture on one end of the spectrum or a single video stream from an ordinary camera on the other end) and on the assessment being performed—see [21] for a review. Most often, the desired output is automated GMA classification (see [22, 23] for reviews), with estimated risks of cerebral palsy prediction as the typical clinical outcome. Direct sensing methods (e.g., accelerometers in [16]) or 3D data from marker-based motion capture have been used [24] but our focus is on indirect sensing and video-based approaches in particular (see Silva et al.[23] for a review). The images of infants can be used to directly extract features that are used for subsequent analyzes and assessment (e.g., [25] for GMA or [26] to analyze U-shaped developmental changes). With the rapid advent of pose estimation methods in computer vision and machine learning, methods that use 2D keypoint extraction in infant recordings are becoming increasingly popular. McCay et al. [27] extracted the positions of infant body keypoints using OpenPose [28], computed histograms of joint orientation and displacement in 2D and then used deep learning to classify abnormal movements (GMA labels). Chambers et al. [29] used the keypoints obtained from OpenPose as input to extract movement features and then applied a Naïve Bayesian Surprise Metric to predict neuromotor development risk. Reich et al. [30] used the positions of the 25 keypoints obtained from OpenPose directly for classification and evaluated the agreement with GMA. Shin et al. [20] obtained the positions of 2D keypoints using AlphaPose [31], estimated joint angle values from the 2D skeletons, analyzed complexity and then applied deep learning for classification (reporting correlations with HINE). Similarly, the risk of autism spectrum disorder (ASD) can be predicted from infant video recordings using input images directly [32] or using 2D pose estimation (OpenPose) first [33].
@@
Methods for automatic human keypoint extraction (for example, eyes, wrists, hips, feet, etc.), also called human pose estimation, from images and videos are rapidly evolving, with performances increasing every year (see [34, 35, 36] for recent surveys). These provide key enabling technology to make automatic motion extraction and analysis from “in the wild” recordings possible. These methods were developed primarily to estimate posture and movement of adult bodies (see [37] for a recent comparison of OpenPose and marker-based motion capture). So far, these methods have been applied to infant videos “as is”. However, the morphology of the infant body is different from the body proportions of the adult, especially in early infancy [38, 39]. Additionally, small infants are typically in supine position (on their back) and move differently than standing adults, which constitutes a dataset which is out of the distribution the adult pose estimation models were trained on. Although automated pose estimation seems to be generally good (see Fig. 1 a-c), in a previous study [11], some limitations of current pose estimation methods in infants were highlighted. For example, OpenPose [28] was found to struggle with some camera angles, body postures, and keypoints. In particular, complex leg positions, such as when legs are crossed in supine position, are not often estimated with high accuracy. This can be seen in Fig. 1 (d): OpenPose misses the left ankle keypoint and places the left knee keypoint on the thigh, very close to the left hip. In the current study, comparing different pose estimation methods, we found similar examples. Fig. 1 (e) shows that HRNet Bottom-Up fails on a complex leg position and fails to place the left shoulder and elbow keypoints, placing them on the right; while (b) shows that HRNet Top-Down hallucinates a second person in the image (see the additional bounding box at the infant’s right forearm), and (f) shows ViTPose not detecting the infant while sitting on the lap of a parent (the image was cropped to not show the parent’s face in the figure).
@@
Figure 1: Examples of successful and erroneous keypoint estimations obtained from the different human pose estimation methods: a) HRnet Bottom-Up (”supine” dataset, infant AA, 17 weeks) successful, b) HRnet Top-Down (”supine” dataset, TH, 10 weeks) with a successful estimation but a second hallucinated person, c) ViTPose (SyRIP image 293) successful, d) OpenPose missing the left ankle keypoint (synthetic infant 1); e) HRNet Bottom-Up failing to estimate a complex leg position (synthetic infant 10); f) ViTPose detecting the parent but not the infant (”Lap” dataset).
Thus, our first goal and contribution is to assess the performance of 2D human pose estimation methods when applied specifically to infants in supine position. In order to provide a more general outlook, we additionally present results on images of infants in a wider variety of postures, including standing or turning away from the camera, as well as with infants on the lap, or sitting or standing in front, of their parent. Only few studies have partially addressed this problem. However, if present, “in the wild” recordings did not constitute the majority of the datasets. Needham et al. [19] used a multi-camera setup in the lab and compared OpenPose, AlphaPose and DeepLabCut. Groos et al. [40] created a large proprietary dataset of human-annotated infant images from both controlled and standardized environments and “in the wild” recordings, and used them to retrain several models, including OpenPose. Their best model achieves on average 81.11% of its estimates being located within an error of 10% of the length of the infant’s face, while the “vanilla” OpenPose had 49.66% of its estimates within this margin of error. Sermpon et al. [41] have found weak correlations between pose estimation features derived from MediaPipe’s estimates and manually annotated movements towards the midline. The preprint of Jahn et al. [42] compared OpenPose, MediaPipe, HRNet, AggPose and VitPose, as well as fine-tuned versions of ViTPose and HRNet (AGMA-HRNet48) with infant recordings in a laboratory context made to specifically target GMA settings (infants in supine position, seen from above and at a diagonal angle). Their results show that ViTPose-huge has the second best results (higher than AGMA-HRNet48, the HRNet model fine-tuned with infant images from a similar GMA context), locating 84.6% of its estimates within 2.5cm of the ground truth, while their re-trained ViTPose locates 93.89% of estimates within this distance.
@@
In this article, we compare the performance of seven methods: Alphapose [31], DeepLabCut / DeeperCut (v2.2.1.1) [43], Detectron2 (v0.6) [44], MediaPipe / BlazePose (v0.10.14) [45, 46], HRNet (MMPose v0.28.0) [47, 48] (Bottom-Up (HRnet BU) and Top-Down (HRnet TD)), OpenPose [28] and ViTPose (MMPose v1.1.0) [49].
@@
The third contribution of this work constitutes in making all the pose estimation methods and tools for the results analysis available to the community. Although the pose estimation algorithms themselves are publicly accessible in their corresponding repositories, we make available the versions we used including the complete environment in public Docker containers (at https://hub.docker.com/u/humanoidsctu) with instructions to simplify their use. In addition, we also share the evaluation scripts and the detailed results (at https://osf.io/x465b/).
@@ Methods
To ensure that comparisons are as fair as possible, we used versions of the pose estimation methods that were trained with the same dataset: COCO [50], although some methods only have available weights that used other datasets in their training, instead of or in addition to COCO (see each method’s respective subsection below or the summary Tab. 1). The methods provide 17 estimated keypoints, with the exception of DeepLabCut, MediaPipe and OpenPose that provide 14, 33 and 18 keypoints respectively. We will call a single set of estimated keypoints a detection. The details of the parameters with which the methods were run are provided in the supplemental materials. All methods except MediaPipe provide a form of confidence values as their own internal evaluation of the estimated result, with higher values indicating higher confidence in the quality of the estimates. This value can be provided for the whole set of keypoints or for individual keypoints (or both). Following the trend in the documentation and methods’ outputs, we will use score when referring to the whole detection value and confidence when referring to the individual keypoints values. The methods follow either one of the two approaches: Top-Down, which first uses a detector to delimit areas of interest in the image that encompass a person to detect and then estimates a set of keypoints within this space; and Bottom-Up, which finds keypoints or sets of keypoints separately on the whole image before joining them into full keypoint sets and separating them into several individuals if necessary.
@@
DeepLabCut/DeeperCut We use the DeepLabCut environment [51], with its human pose estimation DeeperCut [43] based on an R-CNN architecture. It follows a Bottom-Up approach. It has been trained on the MPII dataset [52] and provides 14 keypoints. It provides confidences, but no scores.
@@
MediaPipe/BlazePose While commonly known as MediaPipe [45], the human pose estimation of this environment uses BlazePose [46] and is based on a CNN architecture, following a Top-Down approach. It has been trained on a custom dataset with additional face and hand annotations on top of COCO’s regular 17 keypoints. It provides neither scores nor confidences.
HRnet Bottom-Up and Top-Down [47, 48] are named after the name of the neural architecture used, a High Resolution Network, itself based on CNNs. It has been trained using both COCO and MPII datasets. We use its implementation through the MMPose environment [53], which proposes bottom-up (BU) and top-down (TD) versions of HRNet. HRNet BU provides scores, while HRNet TD provides confidences.